Circular No. 8935

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2491 CYGNI D. K. Lynch, R. W. Russell, and R. J. Rudy, The Aerospace Cor-

poration; C. E. Woodward, University of Minnesota; and G. J. Schwarz, West Chester University, report on SpeX observations (wavelength range 0.8–2.5 μm) of V2491 Cyg (cf. IAUC 8934) obtained at the Infrared Telescope Facility on Apr. 12.56 UT: “The object is definitely a nova, but its spectrum is highly unusual. It has extremely broad lines (FWHM of 5500 km/s) with complex profiles. Fe II may be present in the infrared spectrum, but it is difficult to confirm this object as an ‘Fe II’ nova due to masking of the infrared Fe II emission lines by stronger features. The neutral helium line at 1.083 μm has already appeared, which is early for an ‘Fe II’ nova. This, together with the breadth of the lines, plus strong emission features of N I and N II, suggests that this may be a rarer helium-nitrogen nova. The O I lines indicate a reddening of only E(B-V) = 0.3, a result supported by the blue nature of the overall spectrum. The probable identification of V2491 Cyg with a known x-ray source (Ibarra and Kuulkers,

http://www.astronomerstelegram.org/?read=1473) makes this a poten-

tially important source. Observations at all wavelengths are encouraged.”

Visual magnitude estimates, supplied in part by E. Waagen, AAVSO: Apr. 12.153 UT, 8.4 (P. Schmeer, Bischmisheim, Germany); 12.979, 8.3 (W. Vollmann, Vienna, Austria); 13.372, 9.0 (R. Tyson, Freeport, NY, U.S.A.); 14.436, 9.0 (T. Bretl, Plymouth, MN, U.S.A.).

NR TRIANGULI AUSTRALIS This nova (cf. IAUC 8931) has brightened, as indicated by the following

CCD V magnitudes from G. Bianciardi, University of Siena (remotely using a 0.30-m reflector near Sydney, N.S.W.): Apr. 7.678 UT, 9.22; 8.500, 9.39; 9.699, 9.09; 10.648, 9.20; 11.670, 8.84; 13.712, 8.39.

COMET C/2006 B7 (ODAS) Further to IAUC 8929, E. J. Christensen had sent a potential single-

night recovery of this comet from Mt. Lemmon 1.5-m telescope images on 2005 Oct. 7, noting the comet then to be extremely diffuse with a roughly round 10′′ coma and of magnitude 20.1 in targeted 120-s survey images. These observations were used in the improved orbital elements by G. V. Williams on MPEC 2008-G10, which utilize the A_{3} nongravitational-force parameter normal to the orbit plane, following a demonstration of the value of A_{3} in this particular case by S. Nakano, Sumoto, Japan.

2008 April 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green